OpenStudy (anonymous):

directions=determine the best method to solve each system of equations. then solve the system Question=the owners of the river view restaurant have hired enough servers to handle 17 tables of costumers, and the fire marshal has approved the restaurant for a limit of 56 customers. How many two-seat tables and how many four-seat tables should the owners purchase

OpenStudy (anonymous):

So can you come up with a system of equations first of all?

OpenStudy (anonymous):

The first step is to come up with some variables.

OpenStudy (anonymous):

how do u do that

OpenStudy (anonymous):

"the owners of the river view restaurant have hired enough servers to handle 17 tables of costumers" When you read this, you should think, we need a variable for tables and a variable for customers.

OpenStudy (anonymous):

So we'll say the number of tables is \(t\) and the number of customers is \(c\). Does that make sense so far?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

so what would that part of the equation look like

OpenStudy (anonymous):

"the owners of the river view restaurant have hired enough servers to handle 17 tables of costumers, and the fire marshal has approved the restaurant for a limit of 56 customers." This can be written as:\[ t\le 17\quad c\le 56 \]

OpenStudy (anonymous):

"How many two-seat tables and how many four-seat tables should the owners purchase" Okay now we need a variable for these...

OpenStudy (anonymous):

the answer has to be in the form of the points on a graph and it ether has to be subtraction substitution addition or multiplication

OpenStudy (anonymous):

We can let \(x\) be two-seat tables. We can let \(y\) be four-seat tables.

OpenStudy (anonymous):

We know that \[ x+y = t \]The number of tables is equal to the total of two-seat and four-seat tables. We also know: \[ 2x+4y=c \]The number of customers is 2 for each two-seat and 4 for each four-seat.

OpenStudy (anonymous):

Does this make sense?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Well simplify things a bit, and our final system will be: \[ x+y=17\\ 2x+4y=52 \]

OpenStudy (anonymous):

So we have a system, we just need to solve for it. Which method do you think is best?

OpenStudy (anonymous):

subtraction and it would be 2x+4y=56 because there can be 56 customers

OpenStudy (anonymous):

Oh, ok.

OpenStudy (anonymous):

Are we done then?

OpenStudy (anonymous):

yea thank you